OpenStudy (anonymous):

find the equation of a circle with center at (-1,3) and passes through the point (2,-4)

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

shall i find their distance??

OpenStudy (campbell_st):

find the distance from the centre to the point, that will be the radius... then the standard form is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre... so substitute your values

OpenStudy (anonymous):

the distance gives you the radius.

OpenStudy (anonymous):

\[\sqrt{58}\]

OpenStudy (anonymous):

in this case you already have h and k

OpenStudy (anonymous):

(x+1)^2 +(y-3)^2=58

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

you have to calculate distance from point (2,-4) to (-1,3) by distance formula.....that will be your radius... substitute in \[(x-h)^2 +(y-k)^2=r^2\], where (h,k)= (-1,3)